# Simulation to Rules: A Dual-VLM Framework for Formal Visual Planning

### v1

@@ Abstract
Vision Language Models (VLMs) show strong potential for visual planning but struggle with precise spatial and long-horizon reasoning. In contrast, Planning Domain Definition Language (PDDL) planners excel at long-horizon formal planning, but cannot interpret visual inputs. Recent works combine these complementary advantages by enabling VLMs to turn visual planning problems into PDDL files for formal planning. However, while VLMs can generate PDDL problem files satisfactorily, they struggle to accurately generate the PDDL domain files, which describe all the planning rules. As a result, prior methods rely on human experts to predefine domain files or on constant environment access for refinement. We propose VLMFP, a Dual-VLM-guided framework that can autonomously generate both PDDL problem and domain files for formal visual planning. VLMFP introduces two VLMs to ensure reliable PDDL file generation: A SimVLM that simulates action consequences based on input rule descriptions, and a GenVLM that generates and iteratively refines PDDL files by comparing the PDDL and SimVLM execution results. VLMFP unleashes multiple levels of generalizability: The same generated PDDL domain file works for all the different instances under the same problem, and VLMs generalize to different problems with varied appearances and rules. We evaluate VLMFP with 6 grid-world domains and test its generalization to unseen instances, appearance, and game rules. On average, SimVLM accurately describes 95.5%, 82.6% of scenarios, simulates 85.5%, 87.8% of action sequence, and judges 82.4%, 85.6% goal reaching for seen and unseen appearances, respectively. With the guidance of SimVLM, VLMFP can generate PDDL files to reach 70.0%, 54.1% valid plans for unseen instances in seen and unseen appearances, respectively. Project page: https://sites.google.com/view/vlmfp.
@@ 1 Introduction
In this paper, we address the challenge of visual long-horizon planning by introducing a novel framework, VLM-Guided Formal Planning (VLMFP, illustrated in Fig. 1), a Dual-VLM-guided framework that autonomously generates both problem and domain PDDL files for visual planning. VLMFP integrates two specialized VLMs: a fine-tuned SimVLM that perceives the scenario from visual inputs and simulates action outcomes, and a large GenVLM that generates and iteratively refines PDDL files by aligning their execution with SimVLM’s simulations. Generating both problem and domain PDDL from visual inputs requires object recognition, spatial understanding, reasoning, and PDDL knowledge. We fine-tune a small VLM as SimVLM to strengthen its spatial reasoning, while using a large model as GenVLM for general reasoning and extensive PDDL knowledge.
Importantly, VLMFP achieves multiple levels of generalizability. A generated domain PDDL can be reused across all instances of the same domain, while problem PDDL files can be adapted efficiently for new instances as in-context examples. The framework also transfers well to unseen appearances and even altered environment rules. We evaluate VLMFP on six grid world domains, showing that SimVLM reliably describes scenarios, simulates actions, and determines goal achievement, while GenVLM, guided by SimVLM feedback, generates valid PDDL files that enable planners to solve both seen and unseen problems.
@@ 2 Related Works
LLM and VLM Planning Although Large Language Models (LLMs) have achieved impressive performance on text-based planning tasks (Wei et al., 2022; Yao et al., 2022; Raman et al., 2022; Yao et al., 2024), many real-world planning problems are inherently visual. Relying solely on carefully constructed text inputs is often impractical and cannot capture the rich spatial details in visual environments. This limitation has motivated a growing interest in VLM-based planning, where visual observations provide a more direct and intuitive foundation for reasoning (Driess et al., 2023; Huang et al., 2023; Zhang et al., 2023; Nasiriany et al., 2024). However, current VLMs still fall short in precise spatial understanding (Wu et al., 2024), which restricts their effectiveness on complex planning tasks involving multiple objects and intricate spatial relationships. Some recent works explore Visual Chain-of-Thought (Li et al., 2025; Zhao et al., 2025), which use images as part of the reasoning process while generating the plan. However, this process introduces computational overhead and is hard to apply to long-horizon planning tasks. In this work, we fine-tune SimVLM to enhance its visual–spatial understanding and reasoning, enabling it to guide the generation of PDDL files for solving long-horizon planning problems.
LLM and VLM + PDDL Since existing LLMs lack the ability to reliably perform long-horizon reasoning in complex tasks (Achiam et al., 2023a; Valmeekam et al., 2022; 2023; Kambhampati et al., 2024), recent works have explored combining LLMs with external solvers to augment their reasoning and planning capabilities (Wu et al., 2022; He-Yueya et al., 2023; Pan et al., 2023; Ye et al., 2024; Li et al., 2023; 2024; Hao et al., 2024a; b). Among such solvers, combining LLMs with PDDL-based planners is a powerful option (Silver et al., 2022; Liu et al., 2023; Stein & Koller, 2023; Xie et al., 2023; Stein & Koller, 2023; Guan et al., 2023; Oswald et al., 2024; Mahdavi et al., 2024), as PDDL is designed for precise symbolic reasoning over long-horizon problems. However, these works either require predefined PDDL domain files, human corrections, or environment access to provide feedback while generating PDDL files. In addition, to further enable visual interpretation, many works combine VLMs with PDDL planners (Shirai et al., 2024; Dang et al., 2025) by generating PDDL Problem files based on image observation. However, similarly, as generating PDDL domain files is complicated, these works assume access to predefined domain files. In our work, we use a Dual-VLM-guided framework to autonomously generate both PDDL problem and domain files, without human guidance or environment access.
@@ 3.1 Problem Formulation and Challenges
The goal of VLMFP is, given the problem definition pair $(n_{d},i_{p})$, to come up with an action sequence to achieve the goal. Considering the complementary advantages of VLMs in image perception and natural language understanding, and PPDL in formal planning, VLMFP follows the two-step pipeline, where VLMs are adopted to translate problem definition into PDDL files, and then a PDDL planner is invoked to solve the problem.
However, an unresolved challenge of such a pipeline is that the translation into PDDL files using VLMs is an error-prone process. Specifically, PDDL provides a standardized framework for representing automated planning problems, which involves two files: ❶ A PDDL domain file $f_{d}$, which formally encodes the rules and domain descriptions by defining a set of predicates, actions with preconditions and effects; and ❷ a PDDL problem file $f_{p}$, which formally specifies objects, initial state, and goal conditions for a concrete problem instance. $f_{d}$ and $f_{p}$ can be roughly understood as the formal translation of the domain description $n_{d}$ and problem setup image $i_{p}$, respectively. Please refer to Appendix A.7 for example PDDL files.
@@ 3.2 VLMFP Overview
With these comparative advantages, SimVLM is designed to support GenVLM in two ways. First, SimVLM’s generated description of the spatial relationships in the problem setup image $i_{p}$ is fed to GenVLM, therefore compensating the latter’s weakness in understanding spatial relationships. Second, and more importantly, SimVLM can verify the correctness of the PDDL files by comparing simulation results. Specifically, in VLMFP, there are two ways to predict the consequence of an action. The first is directly by SimVLM; the second is by execution based on the generated PDDL files. If the two results disagree, it indicates a greater probability of incorrect PDDL files, so the disagreement is provided to the GenVLM as feedback to refine the PDDL file generation.
Based on these rationales, VLMFP generates PDDL files via the following four steps:
@@ 3.3 VLM Implementation Details
The fine-tuning process enables SimVLM to better align visual observations with spatially grounded reasoning. In particular, SimVLM learns to generate concise natural language narratives of the initial scenario and to simulate the outcomes of action sequences within these domains. By grounding visual inputs in structured textual descriptions and action simulations, SimVLM serves as a reliable intermediate model that bridges raw perception and the formal representation required for PDDL generation. We show examples of open-source large VLMs’ failure to accurately describe the scenario and simulate actions in Appendix A.4.1.
GenVLM. While SimVLM is tailored for perception and structured simulation, the generation of PDDL domain and problem files requires advanced reasoning to capture action dynamics and a precise understanding of PDDL syntax and conventions to ensure valid domain specifications. To address this, we leverage a large VLM API, GPT-4o (Achiam et al., 2023b), referred to as GenVLM, which has broader reasoning capacity and linguistic knowledge needed for reliable PDDL construction. Guided by the scenario descriptions and simulated outcomes provided by SimVLM, GenVLM generates initial PDDL files and iteratively refines them to resolve inconsistencies. In this way, GenVLM complements SimVLM by bridging high-level reasoning with formal symbolic representation.
@@ 4.1 Domains
We finetune SimVLM and test on six grid world domains: Frozenlake, Maze, Sokoban, Package, Printer, and Overcooked (Towers et al., 2024; Silver & Chitnis, 2020; Jin et al., 2023; Wu et al., 2021). For each domain, we collect data across map sizes ranging from 3 to 8 with varying obstacle probabilities, and create 5–6 distinct visual appearances. In total, this yields a dataset of 430k datapoints. By default, we use this dataset for SimVLM fine-tuning, which is used to evaluate framework VLMFP. Task descriptions, visualizations of different scenario appearances, and dataset statistics are included in Appendix A.2,  A.3, and  A.4.3.
@@ 4.2 SimVLM Performance and Appearance Generalization
We fine-tune a Qwen2-VL-7B model as SimVLM on datasets from six grid world domains. Given a domain description, an image, and an action sequence, SimVLM outputs (1) a natural language problem description, (2) step-by-step reasoning and outcomes, and (3) a judgment on goal achievement. Appendix A.4 provides fine-tuning details and sample inputs and outputs. Before integrating SimVLM into VLMFP, we evaluate its performance on these tasks individually. Since each datapoint has a fixed-format output, we measure performance by exact string matching between the model’s predictions and the ground-truth outputs. The four outputs are referred to as Task Description, Execution Reason, Execution Result, and Goal Reach.
@@
Third, errors in the Task Description do not necessarily affect action simulation. While the Task Description output under U drops by 12.9% on average compared to S, the execution simulation outputs remain equally accurate. This suggests that failures in Task Description typically involve only a small subset of objects, which often do not impact the sampled action sequences. For instance, the Sokoban task contains multiple types of objects: box, goal, wall, agent, floor, and the appearances of these objects are all changed. When the appearances of the objects in a domain change, since SimVLM is not familiar with the new look of objects, it could fail to recognize or localize some objects. However, when the misidentified objects are not directly involved in the sampled action sequence, execution reasoning and results remain correct. This separation highlights that SimVLM is robust in reasoning about action dynamics, even when its initial object recognition is imperfect.
@@ 4.3 VLMFP Performance
We evaluate VLMFP on six domains using GPT-4o (Achiam et al., 2023b). Each problem instance comes with a domain description and an image observation, from which VLMFP generates both PDDL problem and domain files. The problem file is then used as an in-context example to prompt GenVLM to generate PDDL problem files for 100 new instances, with the PDDL Planner validating the plans. We repeat this process for 15 different input instances and calculate the average success rate of finding correct plans, that is, the average success rate of 1500 trials. We follow the same procedure for unseen appearances. Appendix A.5 provides implementation details and failure-case analysis.
@@ 4.5 Generalization on Game Rules
In previous sections, we evaluate how SimVLM and VLMFP generalize to unseen problem instances and appearances. Beyond these, we also explore whether the framework can adapt to changes in the underlying game rules. In FrozenLake, we created 15 rule variants by modifying core mechanics, such as ice hole behavior or movement dynamics (see Appendix A.6 for all rule descriptions). We fine-tune a SimVLM model with these variations and test it on 5 unseen rules. We test 100 problem instances to examine if SimVLM can provide correct Execution Reason and Result for the steps involving unseen rules: 1) Rule1: Since ice is wet, stepping on an ice hole causes the agent to step forward two steps in the same direction. 2) Rule2: Ice holes are teleports to pos-2-2; 3) Rule3: If you step on an ice hole, you have to execute the same actions three times to actually execute it; 4) Rule4: Stepping on an ice hole unlocks a lucky rocket, where you can step forward three steps in the same direction; 5) Rule5: If you step on an ice hole, you freeze and your next action is skipped.
@@
The complexity of the five unseen rules increases gradually from Rule1 to Rule5. Rule 1 has the same effects as a training rule (R7 in Appendix A.6), with entirely different wording and explanation. Rule2 changes the teleport destination; the effect is resolved in a single action. Rule3 increases the repeat count from two to three, influencing later actions and requiring multi-step reasoning. Rule4 extends the rocket mechanic to three steps; since the final position is not explicit, it needs extra spatial inference. Results from Rules 1–4 show that SimVLM handles unseen rules in varied formats and altered dynamics with impressive accuracies of 94.2%, 99.0%, 76.1%, and 59.2%.
Rule 5 introduces a completely novel freezing mechanic never observed in training, and here our method fails to generalize. Interestingly, SimVLM correctly states the reasoning as “The agent stepped on ice hole, next action is skipped” but fails to apply it in the next step, yielding 71.1% success in reasoning but 0% in execution. This indicates that while execution on entirely new dynamics remains challenging, SimVLM demonstrates a promising ability to understand and articulate novel rules, suggesting potential to adapt to entirely unseen rules in the future.

### v2

@@ Abstract
Vision Language Models (VLMs) show strong potential for visual planning but struggle with precise spatial and long-horizon reasoning, while Planning Domain Definition Language (PDDL) planners excel at formal long-horizon planning but cannot interpret visual inputs. Recent works combine these complementary advantages by translating visual problems into PDDL. However, while VLMs can generate PDDL problem files satisfactorily, accurately generating PDDL domain files, which encode planning rules, remains challenging and typically requires human expertise or environment interaction. We propose VLMFP, a Dual-VLM-guided framework that autonomously generates both PDDL problem and domain files for formal visual planning. VLMFP combines a SimVLM that simulates action consequences with a GenVLM that generates and iteratively refines PDDL files by aligning symbolic execution with simulated outcomes, enabling multiple levels of generalization across unseen instances, visual appearances, and game rules. We evaluate VLMFP on 6 grid-world domains and demonstrate its generalization capability. On average, SimVLM achieves 87.3% and 86.0% scenario understanding and action simulation for seen and unseen appearances, respectively. With the guidance of SimVLM, VLMFP attains 70.0%, 54.1% planning success on unseen instances in seen and unseen appearances, respectively. We further demonstrate that VLMFP scales to complex long-horizon 3D planning tasks, including multi-robot collaboration and assembly scenarios with partial observability and diverse visual variations. Project page: https://sites.google.com/view/vlmfp.
@@ 1 Introduction
In this paper, we address the challenge of visual long-horizon planning by introducing a novel framework, VLM-Guided Formal Planning (VLMFP, illustrated in Fig. 1), a Dual-VLM-guided framework that autonomously generates both problem and domain PDDL files for visual planning. VLMFP integrates two specialized VLMs: a fine-tuned SimVLM that perceives the scenario from visual inputs and simulates action outcomes, and a large GenVLM that generates and iteratively refines PDDL files by aligning their execution with SimVLM’s simulations. Generating both problem and domain PDDL from visual inputs requires object recognition, spatial understanding, reasoning, and PDDL knowledge. We fine-tune a small VLM as SimVLM to strengthen its spatial reasoning, while using a large model as GenVLM for general reasoning and extensive PDDL knowledge. This formulation is agnostic to environment dimensionality, allowing the same perception-to-symbolic pipeline to extend beyond grid-based settings to visually grounded 3D planning tasks.
Importantly, VLMFP achieves multiple levels of generalizability. A generated domain PDDL can be reused across all instances of the same domain, while problem PDDL files can be adapted efficiently for new instances as in-context examples. The framework also transfers well to unseen appearances and even altered environment rules. We evaluate VLMFP on six grid world domains, showing that SimVLM reliably describes scenarios, simulates actions, and determines goal achievement, while GenVLM, guided by SimVLM feedback, generates valid PDDL files that enable planners to solve both seen and unseen problems. Beyond grid-world benchmarks, we further examine the scalability of VLMFP on complex long-horizon 3D planning tasks, including assembly and multi-robot collaboration scenarios under partial observability.
@@
  • We demonstrate that VLMFP scales beyond grid-world benchmarks to complex long-horizon 3D tasks with partial observability and diverse visual variations, with an average of 86.4% and 79.8% success rates for unseen instances in seen and unseen appearances, respectively.
@@ 2 Related Works
LLM and VLM Planning Although LLMs have achieved impressive performance on text-based planning tasks (Wei et al., 2022; Yao et al., 2022; Raman et al., 2022; Yao et al., 2024), many real-world planning problems are inherently visual. Relying solely on carefully constructed text inputs is often impractical and cannot capture rich spatial details in visual environments. This limitation has motivated a growing interest in VLM-based planning, where visual observations provide a more direct and intuitive foundation for reasoning (Driess et al., 2023; Huang et al., 2023; Zhang et al., 2023; Nasiriany et al., 2024). However, current VLMs still fall short in precise spatial understanding (Wu et al., 2024), restricting their effectiveness on complex planning tasks involving multiple objects and intricate spatial relationships. Recent works explore Visual Chain-of-Thought (Li et al., 2025; Zhao et al., 2025), using images as part of the reasoning process while generating plans. However, this process introduces computational overhead and is hard to apply to long-horizon planning tasks. In this work, we fine-tune SimVLM to enhance its visual–spatial understanding and reasoning, enabling it to guide the generation of PDDL files for solving long-horizon planning problems.
LLM and VLM + PDDL Since existing LLMs lack the ability to reliably perform long-horizon reasoning in complex tasks (Achiam et al., 2023a; Valmeekam et al., 2022; 2023; Kambhampati et al., 2024), recent works have explored combining LLMs with external solvers to augment their reasoning and planning capabilities (Wu et al., 2022; He-Yueya et al., 2023; Pan et al., 2023; Ye et al., 2024; Li et al., 2023; 2024; Hao et al., 2024a; b). Among such solvers, combining LLMs with PDDL-based planners is a powerful option (Silver et al., 2022; Liu et al., 2023; Stein and Koller, 2023; Xie et al., 2023; Stein and Koller, 2023; Guan et al., 2023; Oswald et al., 2024; Mahdavi et al., 2024), as PDDL is designed for precise symbolic reasoning over long-horizon problems. However, these works either require predefined PDDL domain files, human corrections, or environment access to provide feedback while generating PDDL files. In addition, to further enable visual interpretation, many works combine VLMs with PDDL planners (Shirai et al., 2024; Dang et al., 2025) by generating PDDL Problem files based on image observation. However, similarly, as generating PDDL domain files is complicated, these works assume access to predefined domain files. In our work, we use a Dual-VLM-guided framework to autonomously generate both PDDL problem and domain files, without human guidance or environment access.
@@ 3.1 Problem Formulation and Challenges
The goal of VLMFP is, given the problem definition pair $(n_{d},i_{p})$, to come up with an action sequence to achieve the goal. Considering the complementary strengths of VLMs in image perception and natural-language interpretation, and symbolic methods such as PDDL in formal planning, our framework integrates these advantages to transform visual observations into formal specifications that can be solved symbolically.
However, translating visual inputs into PDDL specifications using VLMs is an error-prone process. PDDL provides a standardized framework for representing automated planning problems, consisting of two files: ❶ a PDDL domain file $f_{d}$, which formally encodes rules by defining predicates and actions with preconditions and effects; and ❷ a PDDL problem file $f_{p}$, which specifies objects, the initial state, and the goal for a concrete instance. $f_{d}$ and $f_{p}$ can be viewed as formal translations of the domain description $n_{d}$ and problem setup image $i_{p}$, respectively. See Appendix E for examples.
@@ 3.2 VLMFP Overview
With these complementary strengths, SimVLM supports GenVLM in two key ways. First, it provides a structured description of the spatial relationships in the problem setup image $i_{p}$, compensating for GenVLM’s weaker spatial perception. Second, SimVLM serves as an execution-checking oracle: by comparing its simulated action outcomes with those derived from the generated PDDL model, any discrepancies signal potential errors and are fed back to GenVLM for refinement.
Based on these rationales, VLMFP generates PDDL files via the following four steps (as in Figure 1):
@@ 4.1 VLM Details
The fine-tuning process enables SimVLM to better align visual observations with spatially grounded reasoning. In particular, SimVLM learns to generate concise natural language narratives of the initial scenario and to simulate the outcomes of action sequences within these domains. By grounding visual inputs in structured textual descriptions and action simulations, SimVLM serves as a reliable intermediate model that bridges raw perception and the formal representation required for PDDL generation. We show examples of API-based large VLMs’ failure to accurately describe the scenario and simulate actions in Appendix B.4.
GenVLM. While SimVLM is tailored for perception and structured simulation, generating PDDL domain and problem files requires advanced reasoning to capture action dynamics and a precise understanding of PDDL syntax and conventions to ensure valid domain specifications. To address this, we leverage a large VLM API, GPT-4o (Achiam et al., 2023b), referred to as GenVLM, which has broader reasoning capacity and linguistic knowledge needed for reliable PDDL construction. Guided by the scenario descriptions and simulated outcomes provided by SimVLM, GenVLM generates initial PDDL files and iteratively refines them to resolve inconsistencies. In this way, GenVLM complements SimVLM by bridging high-level reasoning with formal symbolic representation.
@@ 5.1 Domains
We finetune SimVLM and test on six grid world domains: Frozenlake, Maze, Sokoban, Package, Printer, and Overcooked (Towers et al., 2024; Silver and Chitnis, 2020; Jin et al., 2023; Wu et al., 2021). For each domain, we collect data across map sizes ranging from 3 to 8 with varying obstacle probabilities, and create 5–6 distinct visual appearances. In total, this yields a dataset of 430k datapoints. By default, we use this dataset for SimVLM fine-tuning, which is used to evaluate framework VLMFP. We also evaluate on two more complex 3D embodied tasks, MultiRob and Assembly (Feng et al., 2025; Ji et al., 2025), to assess generalization across environment complexity and visual variability. Domain descriptions, visualizations of different scenario appearances, domain complexity analysis, and dataset statistics are included in Appendix A.1,  A.2,   A.3 and  B.5.
@@ 5.2 SimVLM Performance and Appearance Generalization
We fine-tune a Qwen2-VL-7B (Wang et al., 2024) model as SimVLM on datasets from six grid world domains. Given a domain description, an image, and an action sequence, SimVLM outputs (1) a natural language problem description, (2) step-by-step reasoning and outcomes, and (3) a judgment on goal achievement. Appendix B provides fine-tuning details, sample inputs and outputs. Before integrating SimVLM into VLMFP, we evaluate its performance on these tasks individually. Since each datapoint has a fixed-format output, we measure performance by exact string matching between the model’s predictions and the ground-truth outputs. The four outputs are referred to as Task Description, Execution Reason, Execution Result, and Goal Reach.
@@
Third, errors in the Task Description do not necessarily affect action simulation. While the Task Description output under U drops by 12.9% on average compared to S, the execution simulation outputs remain equally accurate. This suggests that failures in Task Description typically involve only a small subset of objects, which often do not impact the sampled action sequences. For instance, the Sokoban task contains multiple object types: box, goal, wall, agent, floor, and the appearances of these objects are all changed. When the appearances of the objects in a domain change, since SimVLM is not familiar with the new look of objects, it could fail to recognize or localize some objects. However, when the misidentified objects are not directly involved in the sampled action sequence, execution reasoning and results remain correct. This separation highlights that SimVLM is robust in reasoning about action dynamics, even when its initial object recognition is imperfect.
@@ 5.3 VLMFP Performance
We evaluate VLMFP on six domains using GPT-4o (Achiam et al., 2023b). Each problem instance comes with a domain description and an image observation, from which VLMFP generates both PDDL problem and domain files. The problem file is then used as an in-context example to prompt GenVLM to generate PDDL problem files for 100 new instances, with the PDDL Planner validating the plans. We repeat this process for 15 different input instances and calculate the average success rate of finding correct plans, that is, the average success rate of 1500 trials. We follow the same procedure for unseen appearances. Appendix C provides all prompts and failure-case analysis.
@@ 5.5 Generalization on Game Rules
Beyond generalization to unseen instances and visual appearances, we further evaluate whether the framework can adapt to changes in underlying environment rules. In FrozenLake, we construct 15 rule variants by modifying core mechanics such as ice-hole behavior and movement dynamics (see Appendix D for rule descriptions). A SimVLM model is fine-tuned with these variations and evaluated on five unseen rules using 100 problem instances. The unseen rules introduce progressively more complex dynamics, ranging from reworded known behaviors and modified transition effects to novel mechanics that influence multi-step execution (Appendix D). We evaluate whether SimVLM can correctly predict execution reasoning and outcomes for actions involving these unseen rules.
@@
Results are summarized in Table 4. SimVLM achieves strong performance on Rules 1–4, reaching accuracies of 94.2%, 99.0%, 76.1%, and 59.2%, respectively, demonstrating robust generalization to altered rule formulations and modified environment dynamics. These results indicate that SimVLM can successfully interpret and apply previously unseen rule variations when their underlying structure remains related to training dynamics. Performance drops on Rule 5, which introduces a completely novel freezing mechanic not observed during training. Although SimVLM correctly explains the rule behavior (i.e., the next action should be skipped) with 71.1% success rate, it fails to apply this constraint during execution, resulting in correct reasoning but unsuccessful action simulation. This highlights a key limitation when adapting to entirely new dynamics, while also suggesting that SimVLM retains partial semantic understanding of unseen rules.
Overall, these results demonstrate that SimVLM exhibits promising generalization to modified and unseen environment rules, supporting the adaptability of VLMFP beyond fixed planning dynamics.
5.6 VLMFP for Complex 3D Planning Tasks
While previous experiments focus on controlled grid-world domains for systematic evaluation, VLMFP is not restricted to such settings and naturally extends to complex 3D continuous environments. We further evaluate our framework on two complex long-horizon 3D tasks: an assembly task (Feng et al., 2025) and a multi-robot collaboration task (Ji et al., 2025), shown in Figure 4. The domains are challenging because (1) the tasks are fully 3D with no natural-language goal, requiring SimVLM to infer spatial relationships and relational dependencies directly from images, and (2) all block attributes (number, color, shape, position, state), table textures and colors, robot and background colors, and dependency orders are randomized with frequent occlusions, producing highly variable, partially observable configurations.
Figure 4: Visualization of MultiRob and Assembly.
We collect 100k datapoints for each domain and finetune Qwen2-VL-7B as our SimVLM(example inputs and outputs are provided in Appendix B.7). Appearance variation is introduced by randomizing robot, table, background, and object colors, lighting conditions, and camera viewpoints in MultiRob, and by varying block color, position, order, and number under different random seeds in Assembly. Unseen appearances use colors sampled from a different distribution together with unseen random seeds. Following our evaluation protocol, we evaluate SimVLM on 1,000 datapoints each from seen (S) and unseen (U) appearance distributions, and report the end-to-end success rates of VLMFP over 1,500 trials for each setting.
Results and Analysis As shown in Table 5, SimVLM achieves consistently strong performance across all four evaluation metrics, with an average of 96.2%, 96.0% for Assembly and 99.5%, 99.5% for MultiRob for seen and unseen appearances, respectively. This strong SimVLM performance greatly supports PDDL files generation, with VLMFP ’s end-to-end performance: the overall planning success rates are high across both seen and unseen settings for both tasks, achieving 92.4%, 85.9% for Assembly and 80.3%, 73.6% for MultiRob, respectively.
Since SimVLM can already reason over initial and goal images to infer object dependencies, as well as multi-view observations to address occlusions, the framework naturally supports further extensions. For example, incorporating temporal image sequences could enable reasoning over scene dynamics and support planning in evolving environments. Extending our framework to more diverse and realistic 3D tasks is an important direction for future work, particularly given the growing availability of large-scale real-world vision and robotics datasets.
Table 5: SimVLM String matching rate (%) and VLMFP Success rate (%) on 2 3D domains.
Domain Task Descrip. Exec. Reason Exec. Result Goal Reach Average VLMFP
S U S U S U S U S U S U
Assembly 86.9 84.4 99.1 99.1 99.2 99.1 99.4 99.3 96.2 95.5 92.4 85.9
MultiRob 99.9 99.5 99.5 99.4 98.6 99.0 100.0 100.0 99.5 99.5 80.3 73.6